If V(a)=300volt, and V(a->b)=50volt. a test charge (s) was placed 6cm away from A, find the following:
1)the amount and type charge S?
2) the distance between A and B?

B--------A--------S
? 6cm
1
Expert's answer
2018-11-19T15:48:09-0500

1) Positive charge.

V(a)=kq/a

q=V(a)a/k=(300)0.06/(9∙〖10〗^9 )=2∙〖10〗^(-9) C=2 nC.

2)

V(b)=V(a)-V(a->b)=kq/(a+x)

300-50=((2∙〖10〗^(-9))(9∙〖10〗^9))/(0.06+x)

x=0.012 m=1.2 cm.